OpenStudy (meilendurcer):

The length of the base of a triangle is 8.5 cm and its height it 5.4 cm. If the area of the triangle is half of the product of its base length and height, the area of the triangle is_______cm^2

OpenStudy (anonymous):

1/2*8.5*5.4=area

OpenStudy (***[isuru]***):

Hi, follow the info u have Area = (1/2) * length of the base * height All the data is given in the problem Just substitute them here and obtain the value i think u can do that sis ....

OpenStudy (anonymous):

area is \[\frac{ 1 }{ 2 } b*h\]

OpenStudy (anonymous):

i will give medal to isuru you give to me and isuru will give to madrockz :D

OpenStudy (***[isuru]***):

I'll give all the credit to u bro.. It will save the time

OpenStudy (anonymous):

:D ;)